OpenStudy (anonymous):

complete the ordered triple 3x+3y+z=1 x-3y+2z=19 8x-2y+3z=34 someone please help

OpenStudy (amistre64):

whats this?

OpenStudy (amistre64):

you should be a pro by now ;)

OpenStudy (anonymous):

there u r, thought so cant figure out what i did wrong

OpenStudy (amistre64):

let me do this up in notepad and show you a better view of it....matrix style :)

OpenStudy (anonymous):

please

OpenStudy (anonymous):

r u still there doesnt usually take you this long

OpenStudy (amistre64):

z =-1324/23 is one of them, if I did it right :)

OpenStudy (amistre64):

does that sound right to you :)

OpenStudy (anonymous):

i have no idea i didnt do the fractions

OpenStudy (amistre64):

y = -5931/207

OpenStudy (amistre64):

its messy, but lets see if it works in the end :)

OpenStudy (anonymous):

i got x,y,z 2,-3,1

OpenStudy (amistre64):

x = 3324/69

OpenStudy (amistre64):

does it work what you got?

OpenStudy (anonymous):

and of course mathlab says its wrong

OpenStudy (anonymous):

dont know i hate fractions so i eliminate them

OpenStudy (amistre64):

:) i dont think you can get around it this time

OpenStudy (anonymous):

what happen to the superhero genious

OpenStudy (anonymous):

can we start from the begining and work the problem

OpenStudy (amistre64):

were gonna half to , i messed it up somewhere ;)

OpenStudy (anonymous):

ok ready when u r

OpenStudy (amistre64):

I got this so far :) and it looks about like what I had the last time...

OpenStudy (anonymous):

its confusing its not at all what i got i was doing it the way we were earlier

OpenStudy (amistre64):

z = 4 :)

OpenStudy (anonymous):

-6x-6y-2z=-2 x-3y+2z=19 add them and u get -5x-9y=17

OpenStudy (amistre64):

yeah, then it gets messy :)

OpenStudy (amistre64):

y=-3, z=4

OpenStudy (amistre64):

x=2 :) lets double check it ;)

OpenStudy (amistre64):

it just formats better in notepad, and I dont get as lost :)

OpenStudy (anonymous):

-9x-9y-3z=-3 8x-2y+3z=34 add them and you get -x-11y=31

OpenStudy (anonymous):

-5x-9y=17 5x+55y=-155 add them together and get 46y=-138 y=-3

OpenStudy (amistre64):

3(2)+3(-3)+(4)=1 6 -9 +4 = 1 10-9=1...true (2)-3(-3)+2(4)=19 2 +9 +8=19 10+9=19.......true 8(2)-2(-3)+3(4)=34 16 +6+12=34 16 +18=34 .......true (x,y,z) (2 ,-3, 4)

OpenStudy (anonymous):

-x-11y=31 -x-11(-3)=31 -x-33=31 -x=-2 x=2

OpenStudy (amistre64):

you got it :) I did the substitution method....I dont get as many errors that way

OpenStudy (amistre64):

common "Z" ;)

OpenStudy (amistre64):

that didnt work as well as it did in my head :)

OpenStudy (amistre64):

you got 2 out of 3.... work the z.. work the z

OpenStudy (anonymous):

3x+3y+2=1 3(2)+3(-3)+z=1 6x+-9y+z=1 -3+z=1 z=4

OpenStudy (amistre64):

yay!!

OpenStudy (anonymous):

my math was off thats how i got z=-1

OpenStudy (amistre64):

lol..... my math was waaaayyyyyy off to begin with on this one :) BAD FINGERS!!!!

OpenStudy (anonymous):

i forgot how to add i guess

OpenStudy (amistre64):

hey, the cavemen lived without it... and we evolved ;)

OpenStudy (anonymous):

yeah i think people made up this stuff and had to make it confussing so no one could tell them it was wrong

OpenStudy (anonymous):

so when the variables change to a, b, c its the same just different letters

OpenStudy (amistre64):

yep.

OpenStudy (amistre64):

it dont matter what you name them, they are all just different variables in the end

OpenStudy (anonymous):

ok so 6a+3b+c=-20 a-3b+2c=-12 17a-2b+3c=-88

OpenStudy (amistre64):

work it out like you did the other one, and let see your moves lol

OpenStudy (anonymous):

-12a-6b-2c=40 a-3b+2c=-12 add them together get -11a-9b=28

OpenStudy (amistre64):

good so far...

OpenStudy (anonymous):

-18a-9b-3c=60 17a-2b+3c=-88 add them get -a-11b=-28

OpenStudy (amistre64):

-28?

OpenStudy (amistre64):

yeah.......

OpenStudy (amistre64):

-11a-9b=28 -a-11b=-28

OpenStudy (anonymous):

-11a-9b=28 11a+121=308 add them get 112b=336 112(3)=336 b=3

OpenStudy (amistre64):

b = 3, I agree

OpenStudy (anonymous):

-a-11b=-28 -a-11(3)=-28 -a-33b=-28 -a=-61 a=61

OpenStudy (anonymous):

is a right?

OpenStudy (amistre64):

-a -33 = -28 -a = -28+33 -a = 5 a = -5

OpenStudy (anonymous):

i got that the first time thought -61 was too high

OpenStudy (anonymous):

6a+3b+c=-20 6(5)+3(3)+c=-20 30+9+c=-20 c=-59

OpenStudy (amistre64):

c looks good :)

OpenStudy (amistre64):

double check your answers by plugging these values in your eqautions and see if they work

OpenStudy (anonymous):

mathlab says no

OpenStudy (amistre64):

not mathlab, the original equations :)

OpenStudy (anonymous):

i put my answers into mathlab it said they were wrong

OpenStudy (amistre64):

do you have the correct equations? no typos?

OpenStudy (anonymous):

a,b,c 5,3,-59

OpenStudy (anonymous):

c has got to b wrong

OpenStudy (amistre64):

6(5)+3(3)+(-59)=-20 30 +9 -59 = -20 39 - 59 = -20 -20 = -20 (5)-3(3)+2(-59)=-12 5 -9 -118 .....yeah :) 17a-2b+3c=-88

OpenStudy (anonymous):

equation 1 is the only won that ckecked out right

OpenStudy (amistre64):

-12a -6b -2c =40 a -3b +2c =-12 ------------------ -11a -9b = 28

OpenStudy (anonymous):

thats what i got so where did i go wrong

OpenStudy (amistre64):

(6a+3b+c=-20)-3 -18a -9b -3c = 60 17a -2b +3c =-88 ------------------ -a -11b =-28 ....good

OpenStudy (amistre64):

(-a -11b =-28 )-11 11a +121b = 308 -11a - 9b = 28 ------------------ 112b = 336 b = 3

OpenStudy (amistre64):

-11a - 9(3) = 28 -11a = 28+27 -11a = 55 a = -5

OpenStudy (amistre64):

-5 -3(3) +c =-12 -5-9 +c = -12 -14 + c = -12 c = -12+14 c = 2

OpenStudy (anonymous):

ok hold on thats way different then what i got

OpenStudy (amistre64):

6a+3b+c =-20 a-3b+2c =-12 17(-5)-2(-3)+3(2) =-88 -85 +6 +6 that aint working either....

OpenStudy (amistre64):

ow my brain hurts.....

OpenStudy (anonymous):

dont quit on me now

OpenStudy (amistre64):

do substitution......less error :)

OpenStudy (amistre64):

doing it on paper :)

OpenStudy (amistre64):

doing it on paper :)

OpenStudy (anonymous):

a=-5

OpenStudy (amistre64):

(a,b,c) (-5,3,1)

OpenStudy (anonymous):

yeah thats what i just got

OpenStudy (anonymous):

forgot to multiply by -1 to get a

OpenStudy (amistre64):

and I had it, but missed that "2"c....which threw me lol

OpenStudy (anonymous):

i still got 8 more and another page to do gonna b a long night for me

OpenStudy (amistre64):

:) post another question on the left if you need to, this ones a bit long :)